OpenStudy (anonymous):

tan{ cos^-1(4/5) + tan^-1 (2/3)} =?

OpenStudy (ajprincess):

Let \(\cos^-1\large\frac{4}{5}\) be \(\alpha\). Then \(\cos\alpha=\large\frac{4}{5}\) k?

OpenStudy (anonymous):

ok

OpenStudy (ajprincess):

 \(\tan\alpha=\large\frac{3}{4}\) Let \(\tan^-1\large\frac{2}{3}\) be \(\beta\). Then \(\tan\beta=\large\frac{2}{3}\) Nw \(\tan(\alpha+\beta)=\large\frac{\tan\alpha+tan\beta}{1-\tan\alpha\tan\beta}\) \( =\LARGE\frac{\frac{3}{4}+\frac{2}{3}}{1-(\frac{3}{4}*\frac{2}{3})}\) Can u do nw?